A piece of nickel is heated to 107.2 °C. It is then submerged into 235.0 g of water at 22.6 °C that is in a coffee cup calorimeter. The final temperature of the mixture is 28.8 °C. What is the mass of the nickel (in nanograms)? Specific heat of nickel is 0.440 J/g°C. The elemental symbol for nickel is _____?(fill in blank)


1
Expert's answer
2021-01-22T06:24:17-0500

Specific heat(C) of H2O = 4,2 J/g°C

"C = E/(m*\Delta T)"

"\Delta T = 28,8°C - 22,6°C = 6,2°C"

E = 4,24,2 J/g°C *6,2°C*235g = 6119,4J

m(of nickel) = 6119,4J/(0,440J/g°C*78,4°C) = 177,4g

177,4g = 177,4*10-9ng

Answer: 177,4*10-9ng, The elemental symbol for nickel is Ni.